# OmniDepth: Bridging Monocular and Stereo Reasoning with Latent Alignment

### v1

@@ Abstract
Monocular and stereo depth estimation offer complementary strengths: monocular methods capture rich contextual priors but lack geometric precision, while stereo approaches leverage epipolar geometry yet struggle with ambiguities such as reflective or textureless surfaces. Despite post-hoc synergies, these paradigms remain largely disjoint in practice. We introduce OmniDepth, a unified framework that bridges both through iterative bidirectional alignment of their latent representations. At its core, a novel cross-attentive alignment mechanism dynamically synchronizes monocular contextual cues with stereo hypothesis representations during stereo reasoning. This mutual alignment resolves stereo ambiguities (e.g., specular surfaces) by injecting monocular structure priors while refining monocular depth with stereo geometry within a single network. Extensive experiments demonstrate state-of-the-art results: OmniDepth reduces zero-shot generalization error by $40\%$ on Middlebury and ETH3D, while addressing longstanding failures on transparent and reflective surfaces. By harmonizing multi-view geometry with monocular context, OmniDepth enables robust 3D perception that transcends modality-specific limitations. Codes available at https://github.com/aeolusguan/OmniDepth.
@@
Figure 1: OmniDepth bridges the gap between monocular geometry reasoning and stereo pixel matching through latent representation alignment, achieving superior depth accuracy and finer details compared to the monocular model (DepthAnythingV2 [54]), while also surpassing the stereo method (NMRF [13]) in handling textureless surfaces and reflective regions with greater robustness.
@@ 1 Introduction
To bridge this gap, we introduce OmniDepth, a unified framework that harmonizes contextual reasoning and geometric matching via iterative latent representation alignment. At the core of our method is a cross-attentive transformer module that dynamically synchronizes monocular contextual features (encoding object semantics and scene structure) with stereo hypothesis volumes (representing correspondence probabilities). This bidirectional exchange enables monocular priors to disambiguate stereo matching under challenging conditions (e.g., specular or textureless surfaces) via learned structural constraints, while stereo geometry injects metric constraints to rectify monocular depth ambiguities. Unlike prior work [27, 1] relying on post-hoc fusion heuristics, OmniDepth embeds alignment directly within stereo cost aggregation, enabling the two paradigms to mutually refine their representations rather than merely complement one another. As shown in Fig. 1, our method not only achieves consistent improvement in tackling reflective surfaces compared to stereo method NMRF [13] but also realizes depth rectification to recover fine details compared to monocular DepthAnythingV2 (DAv2) model [54].
Our framework bridges monocular and stereo reasoning in two stages. In the pre-alignment stage, the monocular branch captures contextual features of the reference stereo image, while the stereo branch prunes the disparity search space using a Disparity Proposal Network (DPN) [13] for efficient cross-modal interaction. Crucially, the latent alignment stage employs our cross-attentive transformer to iteratively align monocular and stereo representations: it first queries contextual features to guide hypothesis aggregation, then updates the monocular representation using aggregated geometric constraints, enabling bidirectional refinement under dual-task supervision. During inference, OmniDepth produces dual outputs: relative depth [37] from refined contextual features and metric disparity from stereo hypothesis representations. Comprehensive experiments demonstrate that OmniDepth achieves state-of-the-art performance across multiple benchmarks [33, 12, 42, 41, 32], notably reducing zero-shot error by $41.3\%$ on Middlebury [41] and $65.8\%$ on ETH3D [42] versus the leading stereo baseline NMRF [13]. By harmonizing monocular context and stereo matching, OmniDepth transcends modality-specific limitations, enabling robust 3D perception with strong generalization.
@@
    OmniDepth, a unified model jointly predicting relative depth and metric disparity via latent alignment, bridging contextual reasoning and geometric matching.
@@ 3 Approach
Figure 2: Overview of the OmniDepth model. (1) Pre-Alignment: (a) Monocular branch extracts contextual features from the reference image; (b) stereo branch prunes disparity hypotheses (40→2) via DPN [13]. (2) Bidirectional Latent Alignment: (c) Iteratively synchronizes monocular and stereo representations through cross-modal interaction (spatial dimensions are flattened for illustration simplicity). Outputs: relative depth (monocular) and metric disparity (stereo).
@@ 3.4 Rationale of Dual-Outputs
Method RAFT-Stereo [29] DLNR [61] Selective-IGEV [47] NMRF [13] Mocha-Stereo [5] OmniDepth OmniDepth-L
@@ 4.1 Implementation Details
We implement OmniDepth using PyTorch on NVIDIA RTX 3090 GPUs. The AdamW optimizer is adopted with a one-cycle learning rate scheduler across all experiments. Data augmentations similar to [29] are performed. The monocular branch employs a frozen ViT-L version of DAv2 [54], preserving its pretrained generalization capability on real-world data. For stereo feature extraction, we employ the $\tt BasicEncoder$ originated from RAFT [45]. Pretraining is conducted on the Scene Flow [32] dataset for 300K iterations with a batch size of 8. Inputs are randomly cropped to $384\times 768$ resolution, and the learning rate follows a cosine one-cycle schedule peaking at 5e-4.
@@ 4.4 Benchmark Performance
Scene Flow. As shown in Tab. 2, our approach achieves SOTA performance on Scene Flow test set with 3.67% BP-1 error—18.44% better than prior art. Additionally, the scalability of our approach is evidenced by OmniDepth-L, a variant with enlarged stereo feature network (details in suppl.), which achieves 3.25% BP-1 error, demonstrating consistent gains through enhanced representation capacity.
@@ 5 Conclusion
This supplementary material provides additional insights and evaluations to support the findings presented in our main paper. In Sec. A, we elaborate on the details of the stereo feature network of OmniDepth-L. Sec. B presents additional ablation studies, including comparing the performance of our framework when using different monocular foundation backbones, and a comparison with a post-hoc monocular-stereo combination strategy. In Sec. C, we report results from the official ETH3D benchmark, accompanied by a quantitative comparison between our refined monocular depth estimates and well-known monocular foundation models.
A OmniDepth-L variant
@@
In the OmniDepth-L architecture, we utilize the first three “blocks” of the ConvNext-Tiny model. These blocks produce hierarchical feature maps at multiple spatial resolutions: 1/4, 1/8, and 1/16 of the input image size. This multi-scale feature extraction enables the model to capture both fine-grained local details (at higher resolutions) and broader contextual information (at lower resolutions), which are critical for high-quality correspondence estimation. To integrate these multi-resolution features effectively, we employ a DPT fusion module. The DPT fuses the features from different scales into a cohesive representation at 1/4 resolution with a channel dimension of 128. Subsequently, this fused feature map is average-pooled to yield a feature representation at 1/8 resolution, which serves as the input for coarse-level inference within our bidirectional alignment pipeline.
This variant demonstrates the flexibility of our framework, as it can seamlessly scale to incorporate more powerful backbones. By replacing $\tt BasicEncoder$ with enlarged stereo feature network, OmniDepth-L serves as a proof-of-concept for adapting our approach to more sophisticated architectures, potentially yielding superior performance in demanding scenarios (e.g., in Tab. 2).

### v2

# BridgeDepth: Bridging Monocular and Stereo Reasoning with Latent Alignment

@@ Abstract
Monocular and stereo depth estimation offer complementary strengths: monocular methods capture rich contextual priors but lack geometric precision, while stereo approaches leverage epipolar geometry yet struggle with ambiguities such as reflective or textureless surfaces. Despite post-hoc synergies, these paradigms remain largely disjoint in practice. We introduce a unified framework that bridges both through iterative bidirectional alignment of their latent representations. At its core, a novel cross-attentive alignment mechanism dynamically synchronizes monocular contextual cues with stereo hypothesis representations during stereo reasoning. This mutual alignment resolves stereo ambiguities (e.g., specular surfaces) by injecting monocular structure priors while refining monocular depth with stereo geometry within a single network. Extensive experiments demonstrate state-of-the-art results: it reduces zero-shot generalization error by $40\%$ on Middlebury and ETH3D, while addressing longstanding failures on transparent and reflective surfaces. By harmonizing multi-view geometry with monocular context, our approach enables robust 3D perception that transcends modality-specific limitations. Codes available at https://github.com/aeolusguan/BridgeDepth.
@@
Figure 1: Our method bridges the gap between monocular geometry reasoning and stereo pixel matching through latent representation alignment, achieving superior depth accuracy and finer details compared to the monocular model (DepthAnythingV2 [54]), while also surpassing the stereo method (NMRF [13]) in handling textureless surfaces and reflective regions with greater robustness.
@@ 1 Introduction
To bridge this gap, we introduce BridgeDepth, a unified framework that harmonizes contextual reasoning and geometric matching via iterative latent representation alignment. At the core of our method is a cross-attentive transformer module that dynamically synchronizes monocular contextual features (encoding object semantics and scene structure) with stereo hypothesis volumes (representing correspondence probabilities). This bidirectional exchange enables monocular priors to disambiguate stereo matching under challenging conditions (e.g., specular or textureless surfaces) via learned structural constraints, while stereo geometry injects metric constraints to rectify monocular depth ambiguities. Unlike prior work [27, 1] relying on post-hoc fusion heuristics, BridgeDepth embeds alignment directly within stereo cost aggregation, enabling the two paradigms to mutually refine their representations rather than merely complement one another. As shown in Fig. 1, our method not only achieves consistent improvement in tackling reflective surfaces compared to stereo method NMRF [13] but also realizes depth rectification to recover fine details compared to monocular DepthAnythingV2 (DAv2) model [54].
Our framework bridges monocular and stereo reasoning in two stages. In the pre-alignment stage, the monocular branch captures contextual features of the reference stereo image, while the stereo branch prunes the disparity search space using a Disparity Proposal Network (DPN) [13] for efficient cross-modal interaction. Crucially, the latent alignment stage employs our cross-attentive transformer to iteratively align monocular and stereo representations: it first queries contextual features to guide hypothesis aggregation, then updates the monocular representation using aggregated geometric constraints, enabling bidirectional refinement under dual-task supervision. During inference, BridgeDepth produces dual outputs: relative depth [37] from refined contextual features and metric disparity from stereo hypothesis representations. Comprehensive experiments demonstrate that our method achieves state-of-the-art performance across multiple benchmarks [33, 12, 42, 41, 32], notably reducing zero-shot error by $41.3\%$ on Middlebury [41] and $65.8\%$ on ETH3D [42] versus the leading stereo baseline NMRF [13]. By harmonizing monocular context and stereo matching, it transcends modality-specific limitations, enabling robust 3D perception with strong generalization.
@@
    BridgeDepth, a unified model jointly predicting relative depth and metric disparity via latent alignment, bridging contextual reasoning and geometric matching.
@@ 3 Approach
Figure 2: Overview of the BridgeDepth model. (1) Pre-Alignment: (a) Monocular branch extracts contextual features from the reference image; (b) stereo branch prunes disparity hypotheses (40→2) via DPN [13]. (2) Bidirectional Latent Alignment: (c) Iteratively synchronizes monocular and stereo representations through cross-modal interaction (spatial dimensions are flattened for illustration simplicity). Outputs: relative depth (monocular) and metric disparity (stereo).
@@ 3.4 Rationale of Dual-Outputs
Method RAFT-Stereo [29] DLNR [61] Selective-IGEV [47] NMRF [13] Mocha-Stereo [5] BridgeDepth BridgeDepth-L
@@ 4.1 Implementation Details
We implement BridgeDepth using PyTorch on NVIDIA RTX 3090 GPUs. The AdamW optimizer is adopted with a one-cycle learning rate scheduler across all experiments. Data augmentations similar to [29] are performed. The monocular branch employs a frozen ViT-L version of DAv2 [54], preserving its pretrained generalization capability on real-world data. For stereo feature extraction, we employ the $\tt BasicEncoder$ originated from RAFT [45]. Pretraining is conducted on the Scene Flow [32] dataset for 300K iterations with a batch size of 8. Inputs are randomly cropped to $384\times 768$ resolution, and the learning rate follows a cosine one-cycle schedule peaking at 5e-4.
@@ 4.4 Benchmark Performance
Scene Flow. As shown in Tab. 2, our approach achieves SOTA performance on Scene Flow test set with 3.67% BP-1 error—18.44% better than prior art. Additionally, the scalability of our approach is evidenced by BridgeDepth-L, a variant with enlarged stereo feature network (details in suppl.), which achieves 3.25% BP-1 error, demonstrating consistent gains through enhanced representation capacity.
@@ 5 Conclusion
This supplementary material provides additional insights and evaluations to support the findings presented in our main paper. In Sec. A, we elaborate on the details of the stereo feature network of BridgeDepth-L. Sec. B presents additional ablation studies, including comparing the performance of our framework when using different monocular foundation backbones, and a comparison with a post-hoc monocular-stereo combination strategy. In Sec. C, we report results from the official ETH3D benchmark, accompanied by a quantitative comparison between our refined monocular depth estimates and well-known monocular foundation models.
A BridgeDepth-L variant
@@
In the BridgeDepth-L architecture, we utilize the first three “blocks” of the ConvNext-Tiny model. These blocks produce hierarchical feature maps at multiple spatial resolutions: 1/4, 1/8, and 1/16 of the input image size. This multi-scale feature extraction enables the model to capture both fine-grained local details (at higher resolutions) and broader contextual information (at lower resolutions), which are critical for high-quality correspondence estimation. To integrate these multi-resolution features effectively, we employ a DPT fusion module. The DPT fuses the features from different scales into a cohesive representation at 1/4 resolution with a channel dimension of 128. Subsequently, this fused feature map is average-pooled to yield a feature representation at 1/8 resolution, which serves as the input for coarse-level inference within our bidirectional alignment pipeline.
This variant demonstrates the flexibility of our framework, as it can seamlessly scale to incorporate more powerful backbones. By replacing $\tt BasicEncoder$ with enlarged stereo feature network, BridgeDepth-L serves as a proof-of-concept for adapting our approach to more sophisticated architectures, potentially yielding superior performance in demanding scenarios (e.g., in Tab. 2).